A linear combination of two wave functions for a system is also a valid wave function. Find the normalization constant for the combination ψ = B(sin nπx L + sin πx L ) of the wave functions for the n = 1 and n = 2 states of a particle in a box L wide


1
Expert's answer
2022-02-15T16:14:40-0500

"\psi=B(\sin\frac{n\pi x}L+\sin \frac{\pi x}L),"

"\int_0^L\psi^2dx=1,"

"n=1:"

"B=\sqrt\frac 1L,"

"n=2:"

"B=\sqrt\frac 1L."